OpenStudy (anonymous):

Solve: 2x2 – 18x + 28 = 0

OpenStudy (anonymous):

i.e. x^2-9x+14=0 =>x=2 or x=7

OpenStudy (anonymous):

I need to show work though so how you get that

OpenStudy (anonymous):

2x^(2)-18x+28=0 2(x^(2)-9x+14)=0 2(x-2)(x-7)=0 (x-2)(x-7)=0 x=2,7